Consider a fictious virus that causes one to become a zombie! There is a test for this disease with the following sensitivity and specificity:

sensitivity : P(+|zombie virus) = 0.99

specificity : P(−|no zombie virus) = 0.9

The virus can be detected before one turns into a zombie, which takes a week or so, but there is no cure.

At any given time, the proportion of people that have contracted the virus is given by P(Zombie) = 0.001.

Suppose that you have been given a positive test for the virus, what is the probability that you are actually infected?

Show work please.
Answer #1

P(Actually infected | positive test for virus)

A = actually infected

B = positive test for virus

P(A|B)

= P(A and B)/P(B)

P(A and B) = P(B|A)P(A) = 0.99 * 0.001

P(B) = P(B|A)P(A) + P(B|A')P(A')

= 0.99 * 0.001 + (1 - 0.9) ( 1 - 0.001)

= 0.10089

hence

required probability = 0.99* 0.001/ 0.10089

= 0.0098126

Please rate